What will be the change in internal energy of a system that absorbs 300 J of heat and performs 500 J of work?



Calculate the change in internal energy, 400 J of work is done on the system and the system losses 225 J of heat energy.

1
Expert's answer
2022-05-02T13:41:32-0400

1) "\Delta E=Q+W=300\ J -500\ J=-200\ J"

2) "\Delta E=Q+W=-225\ J+400\ J=175\ J"